OpenStudy (kittiwitti1):

CALCULUS I | "Evaluate the following integrals by interpreting it in terms of areas"

OpenStudy (kittiwitti1):

\[a)~\int_{-1}^{4}(2x-1)dx\]\[b)~\int_{0}^{4}(2-\sqrt{16-x^2}dx\]

OpenStudy (kittiwitti1):

What exactly does "interpreting it in terms of areas" mean? Should I use left point, right point or midpoint intervals?

OpenStudy (kittiwitti1):

Speaking of intervals... what should I set \(n\) as?

jimthompson5910 (jim_thompson5910):

I think your teacher wants you to break up the integrals into two pieces, then subtract the two areas eg: this basic form \[\Large \int_{a}^{b}(f(x)-g(x))dx = \int_{a}^{b}f(x)dx-\int_{a}^{b}g(x)dx\]

OpenStudy (kittiwitti1):

Ah. I see o:

OpenStudy (kittiwitti1):

So \(g(x)=1\) in this case?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (kittiwitti1):

Okay... I am assuming I should get this when breaking up then.\[\int_{-1}^{4}(2x-1)dx=\int_{-1}^{4}(2x)-\int_{-1}^{4}(1)\]

OpenStudy (kittiwitti1):

I am not exactly sure how to handle the broken-up parts though... should I use \(\Sigma\) summation?

jimthompson5910 (jim_thompson5910):

have you learned about antiderivatives?

OpenStudy (kittiwitti1):

Yes Oops, I forgot the \(dx\) at the end it seems

OpenStudy (kittiwitti1):

Solving an integral \(\rightarrow\) antiderivative of integrated function?

jimthompson5910 (jim_thompson5910):

I would use that method over Riemann Sums

jimthompson5910 (jim_thompson5910):

if dy/dx = 2x then y = ???

OpenStudy (kittiwitti1):

\(y=x^2\) when \(\frac{dy}{dx}=2x\) because \((x^2)'=2x\)

jimthompson5910 (jim_thompson5910):

yes, good

jimthompson5910 (jim_thompson5910):

well +C but that +C will go away after you evaluate the endpoints

OpenStudy (kittiwitti1):

So then...\[\int_{-1}^{4}(2x)dx\rightarrow x^2\]and then...\[\int_{-1}^{4}(1)dx\rightarrow x\]

OpenStudy (kittiwitti1):

I am not sure if I should add a C to those or not...

jimthompson5910 (jim_thompson5910):

no need because the C's will cancel

jimthompson5910 (jim_thompson5910):

once you get to x^2, evaluate it at the endpoints and subtract same for the other piece

OpenStudy (kittiwitti1):

Evaluate at... endpoints?

jimthompson5910 (jim_thompson5910):



OpenStudy (kittiwitti1):

OH. Well, I feel ridiculously stupid now.

jimthompson5910 (jim_thompson5910):